Hendeles and Weinberger further state an infusion of 0.85 mg/kg/hr of theophylline will produce an eventual plasma theophylline concentration of 10 mcg/ml in children 1-9 years old (plasma concentrations, however, should be monitored after 12-24 hours to avoid toxicity).

a) If Hendeles and Weinberger are correct about the Cpss of 10mcg/ml, what would be the K and t1/2 in this example child?

(b) What infusion rate (Ro) of aminophylline should therefore be given to a 3-year-old. 15 kg child to achieve 12 mcg/ml?

(c) When will steady-state be reached (in hours)?
Answer #1

C= Ko/Cl

C= Steady state concentration

Ko=Infusion rate

Cl=Clearance

10 * 10^-3 = 0.85/Cl

Cl= 85 ml/kg/hour

B) C= Ko/Cl

12* 10^-3= Ko/85

Ko= 1.02 mg/kg/hr

C) Steady state is reaches in 4-5 half lives.